OpenStudy (gabbyalicorn):

Solve. –40 = 5x A. –5 B. –8 C. –35 D. –200

pooja195 (pooja195):

Divide both sides by 5

OpenStudy (gabbyalicorn):

what do you mean both sides?

OpenStudy (is3535):

B

pooja195 (pooja195):

divide both sides by 5 5/5=1 right so you are left with x x=40/5=?

OpenStudy (gabbyalicorn):

40 ÷ 5 = 8

pooja195 (pooja195):

*-40/5

pooja195 (pooja195):

So you get -8 x=-8

pooja195 (pooja195):

So you get -8 x=-8

OpenStudy (gabbyalicorn):

ah, right because 40 is the bigger number. Thank you.

pooja195 (pooja195):

:)

OpenStudy (gabbyalicorn):

Also, is that character from an anime? Or is it fan art or something like that.